two buildings are 60 feet apart acrooss a street. a person on top of the building finds the angle of elevation of the roof of the taller building to be 20 degrees and the angke of depression of its base to be 35 degrees. how tall is the taller building to the nearest foot?
Answer by mananth(16946)

Angle of elevation = 20
angle of depression to the foot = 35
.
let h1 be the difference in heights of the building
Let h2 be the height of the smaller building
h1+h2 = H the height of the taller building.
.
tan 20 + tan 35 = h1/D+h2/D
tan 20 + tan 35= h1+h2 /D
d*(tan 20 + tan 35)= H
60*(tan 20 + tan 35)=H
163 feet